OpenStudy (anonymous):

GI bisects with DGH so that m DGI is x-3 and m IGH is 2x-13 . Find the value of x. HELP!

jimthompson5910 (jim_thompson5910):

Since GI bisects angle DGH, this means that angles DGI and IGH are congruent

jimthompson5910 (jim_thompson5910):

so DGI = IGH x-3 =2x-13

OpenStudy (anonymous):

so what does x equal.. lol

jimthompson5910 (jim_thompson5910):

solve for x and tell me what you get

OpenStudy (anonymous):

idk how..

jimthompson5910 (jim_thompson5910):

Start by adding 3 to both sides

OpenStudy (anonymous):

add 3 to what..?

jimthompson5910 (jim_thompson5910):

x-3 =2x-13 x-3+3 =2x-13+3 .. Add 3 to both sides. x = 2x-10

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):

that makes no sence..

jimthompson5910 (jim_thompson5910):

alright, let me find a few pages on solving linear equations

OpenStudy (anonymous):

ugh

jimthompson5910 (jim_thompson5910):

here's one page http://www.purplemath.com/modules/solvelin.htm and here's another http://www.regentsprep.org/Regents/math/ALGEBRA/AE2/LSolvEq.htm let me know if they make sense or not

jimthompson5910 (jim_thompson5910):

the basic idea is to isolate x you do this by moving all the x terms to one side and all the non x terms to the other

OpenStudy (anonymous):

so do you know the answer..

jimthompson5910 (jim_thompson5910):

i do, but I want to help you get it and not just give you the answer

OpenStudy (anonymous):

okay so what do i do now?

jimthompson5910 (jim_thompson5910):

The next step is to subtract 2x from both sides

jimthompson5910 (jim_thompson5910):

x = 2x-10 x-2x = 2x-10 - 2x -x = -10 then what?

OpenStudy (anonymous):

uh.. i have no idea..

OpenStudy (anonymous):

add x? idkk.. /:

jimthompson5910 (jim_thompson5910):

if -x = -22 for example, then x = 22

OpenStudy (anonymous):

so it equals 10?

jimthompson5910 (jim_thompson5910):

you got it, x = 10

OpenStudy (anonymous):

so thats the answer?

jimthompson5910 (jim_thompson5910):

it is

OpenStudy (anonymous):

lol thanks so much c:

jimthompson5910 (jim_thompson5910):

yw